B. Let f(t) be the solution of the initial value problem dy dt + 6y = exp(lt) with y(0) = 1/7. Find f(t). f(t) = [e^(t)j7

Can you explain me how this is right with the steps, please?
Answer #1

* e +63 set you at d{ y+6y} = {et? Cleats > Lly}+6 289} = 5-a 1 . => 16 465) - 571 yol). + 6 483) 3-1 => 763 8+6) - 12 = 1/